OpenStudy (anonymous):

Factor Completely: 18a^4-24a^3b+8a^2b^2 The answer is 2a^2(3a-2b)^2, but I have no idea how to solve it. This is as far as I can get by myself: 18a^4-24a^3b+8a^2b^2 2a^2(9a^2-12ab+4b^2) I don't know how to continue factoring from there. I'd appreciate any help you could give me. :)

OpenStudy (anonymous):

Oh, never mind, I figured it out. To those of you who'd like to know: 18a^4-24a^3b+8a^2b^2 = 2a^2(9a^2-12ab+4b^2) Factor: 9a^2-12ab+4b^2 Product: 9*4=36 --> -6*-6 Sum: -12 --> -6+(-6) Bring down 9a^2, both factors, and then 4b^2 = 9a^2-6ab-6ab+4b^2 Factor by grouping: = (9a^2-6ab)(-6ab+4b^2) = 3a(3a-2b)-2b(3a-2b) = (3a-2b)^2

OpenStudy (anonymous):

excellent work :)

OpenStudy (anonymous):

Thank you! It took me embarrassingly long to solve, though. Lol.

OpenStudy (anonymous):

Its kinda like doing FOIL in reverse